OpenStudy (anonymous):

Find the perimeter of a square with s = 3.5 inches.

OpenStudy (anonymous):

x 4

OpenStudy (anonymous):

3.5 * 4 = 14 inches

OpenStudy (turingtest):

Perimeter of a rectangle=2(length+width) for a square length=width=s so your equation becomes P=2(l+w)=2(s+s)=2(2s)=4s